# Topological consequences of null-geodesic refocusing and applications to $Z^{x}$ manifolds

### v2

@@ 3 Related results about $Y^{x}_{l}$ manifolds and refocusing spacetimes
The construction of Proposition 3.26 shows that there is a positive Legendrian isotopy from a fiber $S_{x}^{\ast}M$ to a fiber $S_{y}^{\ast}M$ of $S^{\ast}M$. Since $S^{\ast}M$ is symmetric with respect to choice of fibers, there also is a positive Legendrian isotopy from $S_{y}^{\ast}M$ to $S_{x}^{\ast}M$. Concatenating these isotopies yields a non-negative Legendrian isotopy from $S^{\ast}_{x}M$ to itself. By [11, Remark 1.1] this proves the claim. ∎
@@
There is a different way to construct a non-negative Legendrian isotopy from one fiber of $S^{\ast}M$ to itself in the case where $M$ is the Cauchy surface of a globally hyperbolic strongly refocusing spacetime $(X,g)$. Let $p,q\in X$ have the same sky. By [14, Proposition 1.1], [1, Proposition 5.8] there is a non-negative Legendrian isotopy between skies of any causally related points. Since $p$ and $q$ are causally related with the same sky, there is a non-negative Legendrian isotopy from $\mathrm{sky}_{p}$ to itself. Fixing some Cauchy surface $M\subset X$ with $p\in M$ and identifying $\mathcal{N}$ with $S^{\ast}M$ we get a non-negative Legendrian isotopy from a fiber of $S^{\ast}M$ to itself.
@@ 4 Observer-Refocusing Spacetimes
Let $X$ be a Lorentzian manifold and $p\in X$. We denote by $\mathcal{L}_{p}$ the set of future-pointing null-directions at $p$, i.e. we define $\mathcal{L}_{p}=\{[v]_{\sim}\mid v\text{ future pointing and lightlike}\}$ where $\sim$ is the equivalence relation on the set of null-vectors defined through $v_{1}\sim v_{2}\iff\exists\lambda>0:v_{1}=\lambda v_{2}$. We call a smooth submanifold $L\subset T_{p}X$ a lightsphere at $p$ if every $v\in L$ is lightlike and if the natural map from $L$ to $\mathcal{L}_{p}$ is a homeomorphism.
Such lightspheres always exist, for example one can take any auxiliary Riemannian metric $h$ on a Lorentzian manifold $(X,g)$ and define $L=\{v\in T_{p}X\mid v\text{ future pointing and lightlike and }h(v,v)=1\}$.
@@
Let $(X.g)$ be a globally hyperbolic spacetime and very weakly refocusing with respect to some $p\in X$. Then the Cauchy surface of $(X,g)$ is compact.
@@
For each $v\in L$ let $t_{v}$ be such that $\alpha_{v}(t_{v})\in I^{+}(p)$. Take some open neighborhood $U_{v}$ around $v$ such that for all $w\in\overline{U_{v}}$ we have that $\alpha_{w}(t_{v})\in I^{+}(p)$. Since $L$ is compact, we can cover it with a finite number of such neighborhoods $U_{v_{1}},\dots,U_{v_{n}}$. Consider the set $K=\bigcup_{i=1}^{n}\{\alpha_{w}(t_{v_{i}})\mid w\in\overline{U_{v_{i}}}\}$, which is compact as a finite union of compact sets. Let $S\subset X$ be a Cauchy surface with $K\subset I^{-}(S)$ and consider the set $J^{+}(p)\cap S$. Obviously $J^{+}(p)\cap S$ is a non-empty, closed subset of $S$. Further, we get that
@@
If $(M,h)$ is a $\tilde{Z}^{x}$-manifold, then $M$ is compact. In particular every $Z^{x}$ manifold is compact.
@@
Let $(X,g)$ be a globally hyperbolic, $p\in X$ and $\gamma$ a timelike curve such that $(X,g)$ is observer-refocusing with respect to $p$ and $\gamma$. If $p=\gamma(0)$, then $(X,g)$ is very weakly refocusing. In particular $(X,g)$ has compact Cauchy surface.
@@
For every future-directed timelike vector $v\in T_{p}X$ there is some $q\in\mathrm{im}(\gamma)$ and $t>0$ such that $\alpha_{v}(t)\in\mathrm{im}(\gamma)$, i.e. $\alpha_{v}(t)=\gamma(s)$ for some $s\in\mathbb{R}$. We have $\alpha_{v}(t)\in J^{+}(p)\setminus\{p\}$. Since $\gamma$ is timelike we must have $s>0$ and hence $\alpha_{v}(t)=\gamma(s)\in I^{+}(p)$. We apply Lemma 4.30 and are done. ∎
@@
Let $(X,g)$ be a globally hyperbolic spacetime, $p\in X$, let $L$ be a lightsphere at $p$, let $A\subseteq L$ be a dense subset of $L$, and let $T>0$. If for every null-geodesic $\alpha$ with $\alpha(0)=p$ and $\alpha^{\prime}(0)\in A$ we have that $\alpha(T)\in I^{+}(p)$, then $X$ has compact Cauchy surface.
@@
We can assume that $(X,g)$ is null-geodesically complete, else consider a globally hyperbolic spacetime conformally equivalent to $(X,g)$ that is. Let $v\in L$ and let $v_{n}$ be a sequence in $A$ converging towards $v$ Let $t_{n}$ be such that for every $n\in\mathbb{N}$ we have that $\alpha_{v_{n}}(t_{n})$ is the first null-cut point of $\alpha|_{[0,t_{n}]}$. By assumption the sequence $(t_{n})_{n\in\mathbb{N}}$ is bounded. Without loss of generality assume that the sequence converges, else go to a subsequence. Since the null-cut locus $C_{N}^{+}(p)$ is closed (see [2]), we have $\alpha_{v}(t)\in C_{N}^{+}(p)$. Therefore $\alpha_{v}(t^{\prime})\in I^{+}(p)$ for all $t^{\prime}>t$. The result follows by applying Lemma 4.30. ∎
@@
Since the result is entirely local, we can assume $M=\mathbb{R}^{n}$ and $x=0$. Without loss of generality, assume that $x_{n}\to x$ and that $x_{n}\neq x$ for all $n$, else take a subsequence. Further, without loss of generality, assume that $\lim_{n\to\infty}\frac{x_{n}}{||x_{n}||}=v$ for some unit vector $v$ and that $||x_{n+1}||<\frac{1}{2^{n}}||x_{n}||$ for all $n$, else we again take a subsequence. Then the sequence $(w_{n})_{n\in\mathbb{N}}$ defined through $w_{n}\coloneqq\frac{x_{n}-x_{n+1}}{||x_{n}-x_{n+1}||}$ converges to $v$. We define the curve $\gamma$ as follows. We connect $x_{n+1}$ to $x_{n}$ by a $\mathcal{C}^{1}$ curve $\gamma_{n}:[0,a_{n}]$, parametrized by arc-length, satisfying $\gamma_{n}(0)=x_{n+1}$, $\gamma_{n}(a_{n})=x_{n}$, $\gamma_{n}^{\prime}(0)=w_{n+1}$ and $\gamma_{n}^{\prime}(a_{n})=w_{n}$. Since the sequence $(w_{n})_{n\in\mathbb{N}}$ converges to $v$, the interpolations can be done in a way (for example by interpolating with cubic splines) such that that $\lim_{n\to\infty}\sup_{t\in[0.a_{n}]}||\gamma_{n}^{\prime}(t)-v||=0$ and that $\sum_{n\in\mathbb{N}}a_{n}<\infty$. We then define $\gamma$ as the infinite concatenation
@@
By construction $\gamma$ is $\mathcal{C}^{1}$, as we always have $\gamma_{n+1}^{\prime}(a_{n+1})=w_{n+1}=\gamma_{n}(0)$. Further we have $\lim_{t\to 0}\gamma(t)=x$ as well as $\lim_{t\to 0}\gamma^{\prime}(t)=v$, so $\gamma$ extends to a $\mathcal{C}^{1}$ curve defined on $[0,\sum_{n\in\mathbb{N}}a_{n}]$ with the desired property.
@@
Let $(X,g)$ be a Lorentzian manifold and let $h:[0,1]\to T_{p}X$ be a curve such that $h(t)$ is lightlike for each $t\in[0,1]$. Then $h^{\prime}(0)$ is not timelike. Hence $d\exp_{0}(h^{\prime}(0))=\frac{d}{dt}|_{t=0}\exp(h(t))$ is not timelike.
@@
We have that $g(h(t),h(t))=0$ for all $t\in[0,1]$. Therefore $02g(h(0),h^{\prime}(0))=\frac{d}{dt}|_{t=0}g(h(t),h(t))=$ 0. This means that $h^{\prime}(0)$ is perpendicular to the null-vector $h(0)$, and hence $h^{\prime}(0)$ cannot be timelike. ∎

### v4

@@ 3 Related results about $Y^{x}_{l}$ manifolds and refocusing spacetimes
The construction of Proposition 3.26 shows that there is a positive Legendrian isotopy from a fiber $S_{x}^{\ast}M$ to a fiber $S_{y}^{\ast}M$ of $S^{\ast}M$. Since $S^{\ast}M$ is symmetric with respect to choice of fibers, there also is a positive Legendrian isotopy from $S_{y}^{\ast}M$ to $S_{x}^{\ast}M$. Concatenating these isotopies yields a non-negative Legendrian isotopy from $S^{\ast}_{x}M$ to itself. By [12, Remark 1.1] this proves the claim. ∎
@@
There is a different way to construct a non-negative Legendrian isotopy from one fiber of $S^{\ast}M$ to itself in the case where $M$ is the Cauchy surface of a globally hyperbolic strongly refocusing spacetime $(X,g)$. Let $p,q\in X$ have the same sky. By [15, Proposition 1.1], [1, Proposition 5.8] there is a non-negative Legendrian isotopy between skies of any causally related points. Since $p$ and $q$ are causally related with the same sky, there is a non-negative Legendrian isotopy from $\mathrm{sky}_{p}$ to itself. Fixing some Cauchy surface $M\subset X$ with $p\in M$ and identifying $\mathcal{N}$ with $S^{\ast}M$ we get a non-negative Legendrian isotopy from a fiber of $S^{\ast}M$ to itself.
@@ 4 Observer-Refocusing Spacetimes
Let $X$ be a Lorentzian manifold and $p\in X$. We denote by $\mathcal{L}_{p}$ the set of future-pointing null-directions at $p$, i.e. we define $\mathcal{L}_{p}=\{[v]_{\sim}\mid v\text{ future-pointing and null}\}$ where $\sim$ is the equivalence relation on the set of null-vectors defined through $v_{1}\sim v_{2}\iff\exists\lambda>0:v_{1}=\lambda v_{2}$. Let $\mathcal{N}_{p}X$ be the space of null-vectors in $T_{p}X$. We call a smooth submanifold $L\subset\mathcal{N}_{p}X$ a lightsphere at $p$ if the natural map from $L$ to $\mathcal{L}_{p}$ is a homeomorphism.
Such lightspheres always exist, for example one can take any auxiliary Riemannian metric $h$ on a Lorentzian manifold $(X,g)$ and define $L=\{v\in T_{p}X\mid v\text{ future-pointing and null and }h(v,v)=1\}$.
@@
Let $(X,g)$ be a globally hyperbolic spacetime which is very weakly refocusing with respect to some $p\in X$. Then the Cauchy surface of $(X,g)$ is compact.
@@
Let $L$ be some lightsphere at $p$. For each $v\in L$ let $t_{v}$ be such that $\alpha_{v}(t_{v})\in I^{+}(p)$. Take some open neighborhood $U_{v}$ around $v$ such that for all $w\in\overline{U_{v}}$ we have that $\alpha_{w}(t_{v})\in I^{+}(p)$. Since $L$ is compact, we can cover it with a finite number of such neighborhoods $U_{v_{1}},\dots,U_{v_{n}}$. Consider the set $K=\bigcup_{i=1}^{n}\{\alpha_{w}(t_{v_{i}})\mid w\in\overline{U_{v_{i}}}\}$, which is compact as a finite union of compact sets. Let $S\subset X$ be a Cauchy surface with $K\subset I^{-}(S)$ and consider the set $J^{+}(p)\cap S$. Obviously $J^{+}(p)\cap S$ is a non-empty, closed subset of $S$. Further, we get that
@@
If $(M,h)$ is a $\tilde{Z}^{x}$ manifold, then $M$ is compact. In particular every $Z^{x}$ manifold is compact.
@@
Let $(X,g)$ be a globally hyperbolic spacetime, $p\in X$ and $\gamma$ a timelike curve such that $(X,g)$ is observer-refocusing with respect to $p$ and $\gamma$. If $p=\gamma(0)$, then $(X,g)$ is very weakly refocusing. In particular $(X,g)$ has a compact Cauchy surface.
@@
For every future-directed null vector $v\in T_{p}X$ there is some $q\in\mathrm{im}(\gamma)$ and $t>0$ such that $\alpha_{v}(t)\in\mathrm{im}(\gamma)$, i.e. $\alpha_{v}(t)=\gamma(s)$ for some $s\in\mathbb{R}$. We have $\alpha_{v}(t)\in J^{+}(p)\setminus\{p\}$. Since $\gamma$ is timelike we must have $s>0$ and hence $\alpha_{v}(t)=\gamma(s)\in I^{+}(p)$. We apply Lemma 4.30 and are done. ∎
4.33 Remark.
In general, the spacetimes $(X,g)$ we consider will not be null-geodesically complete, so expressions of the type $\alpha_{v}(t)$ might become undefined for large $t$. However, every globally hyperbolic spacetime $(X,g)$ is conformally equivalent to a spacetime $(X^{\prime},g^{\prime})$ which is null-geodesically complete (see [17], where this is proven for all strongly causal spacetimes). Our proofs rely only on the causal structure and topology of our spacetime, both of which are preserved by conformal isomorphisms. So, when convenient, we will assume without loss of generality that our spacetime $(X,g)$ is null-geodesically complete.
@@
Let $(X,g)$ be a globally hyperbolic spacetime, $p\in X$, let $L$ be a lightsphere at $p$, let $A\subseteq L$ be a dense subset of $L$, and let $T>0$. If for every null-geodesic $\alpha$ with $\alpha(0)=p$ and $\alpha^{\prime}(0)\in A$ we have that $\alpha(T)\in I^{+}(p)$, then $X$ has a compact Cauchy surface.
@@
By Remark 4.33, we can assume that $(X,g)$ is null-geodesically complete. Let $v\in L$ and let $v_{n}$ be a sequence in $A$ converging towards $v$. Let $t_{n}$ be such that for every $n\in\mathbb{N}$ we have that $\alpha_{v_{n}}(t_{n})$ is the null-cut point of $\alpha_{v_{n}}|_{[0,t_{n}]}$. By assumption the sequence $(t_{n})_{n\in\mathbb{N}}$ is bounded by $T$. Without loss of generality assume that the sequence converges to some value $t$, else go to a subsequence. Since the null-cut locus $C_{N}^{+}(p)$ is closed (see [2, Proposition 6.5]), we have $\alpha_{v}(t)\in C_{N}^{+}(p)$. Therefore $\alpha_{v}(t^{\prime})\in I^{+}(p)$ for all $t^{\prime}>t$. The result follows by applying Lemma 4.30. ∎
4.35 Lemma.
Let $(X,g)$ be a globally hyperbolic spacetime of dimension at least $3$, let $p\in X$, let $L$ be some lightsphere at $p$ and let $T>0$. If $L$ can be written as a countable union of closed sets
$L=\bigcup_{i\in I}K_{i}$
such that for each $i\in I$ and each $v\in\mathrm{int}(K_{i})$ we have $\alpha_{v}(T)\in I^{+}(p)$, then the Cauchy surface of $(X,g)$ is compact.
Proof.
By Remark 4.33, we can assume that $(X,g)$ is null-geodesically complete. Each set $K_{i}$ is closed in $L$, hence the boundaries $\partial_{L}(K_{i})$ are nowhere dense in $L$. Let $A=L\setminus\bigcup_{i\in I}\partial_{L}(K_{i})$, then $A$ is the complement of a countable union of nowhere dense sets, hence $A$ is dense by the Baire category theorem. Further, if $v\in A$, then there is some $i\in I$ with $v\in K_{i}\setminus\partial_{L}(K_{i})=\mathrm{int}_{L}(K_{i})$, and hence $\alpha_{v}(t)\in I^{+}(p)$ for all $t\geq T$. Since $A$ is dense, $X$ has a compact Cauchy surface by Lemma 4.34.
@@
[…]
@@
By construction $\gamma$ is $\mathcal{C}^{1}$, as we always have $\gamma_{n+1}^{\prime}(a_{n+1})=w_{n+1}=\gamma_{n}^{\prime}(0)$. Further we have $\lim_{t\to 0}\gamma(t)=x$ as well as $\lim_{t\to 0}\gamma^{\prime}(t)=v$, so $\gamma$ extends to a $\mathcal{C}^{1}$ curve defined on $0,\sum_{n\in\mathbb{N}}a_{n}]$ with the desired property.
@@
Let $(X,g)$ be a Lorentzian manifold and let $h:[0,1]\to T_{p}X$ be a curve such that $h(t)$ is null for each $t\in[0,1]$. Then $d\exp_{h(0)}(h^{\prime}(0))=\frac{d}{dt}|_{t=0}\exp(h(t))$ is not timelike.
@@
We have that $g(h(t),h(t))=0$ for all $t\in[0,1]$. Therefore $2g(h(0),h^{\prime}(0))=\frac{d}{dt}|_{t=0}g(h(t),h(t))=0$. This means that $h^{\prime}(0)$ is perpendicular to the null-vector $h(0)$. Let us denote $v=h(0)$ and $w=h^{\prime}(0)$. By the Gauss lemma,
$g(d\exp_{v}(v),d\exp_{v}(w))=g(v,w)=0.$
Here $d\exp_{v}(v)$ is a null-vector since $v$ is one. Therefore $d\exp_{v}(w)$ is orthogonal to a null-vector and hence cannot be timelike. ∎
@@
4.39 Lemma.
Let $(X,g)$ be a globally hyperbolic spacetime of dimension at least $3$, let $p\in X$ and let $\gamma$ be some timelike curve in $X$. Let $L$ be a lightsphere at $p$ and $v\in L$ and let $T>0$. Let $A=\{v\in L\mid\alpha_{v}|_{(0,T]}\text{ intersects }\gamma\}$ and $B=\{v\in L\mid p\text{ has a conjugate point }q\in\mathrm{im}(\gamma)\text{ along }\alpha_{v}\})$. Then $\mathrm{int}(A)\subseteq B$.
Proof.
By Remark 4.33, we can assume that $(X,g)$ is null-geodesically complete. Let $v\in\mathrm{int}(A)$. Since $\dim(X)\geq 3$ implies $\dim(L)\geq 1$, there is some sequence $(v_{n})_{n\in\mathbb{N}}$ in $A$ converging to $v$. For each $n\in\mathbb{N}$ let $t_{n}\in(0,T])$ and $q_{n}\in\mathrm{im}(\gamma)$ be such that $\alpha_{v_{n}}(t_{n})=q_{n}$. Since all $q_{n}$ are contained in the compact set $\{\exp(tv)\ \mid v\in L,t\in[0,T]\}$, we can assume that $(t_{n})_{n\in\mathbb{N}}$ and $(q_{n})_{n\in\mathbb{N}}$ converge to some $t$ and $q$ respectively, else go to convergent subsequences. By Lemma 4.36 there exists a $\mathcal{C}^{1}$ curve $\beta:[0,1]\to\mathcal{N}_{p}X$ with $\beta(0)=tv$ and $\beta^{\prime}(0)\neq 0$ and such that for every $\varepsilon>0$, there are infinitely many $t_{n}v_{n}\in\beta([0,\varepsilon])$. Consider the $\mathcal{C}^{1}$ curve $\alpha:[0,1]\to X$ defined through $\alpha(s)\coloneqq\exp(\beta(s))$. Then for every $\varepsilon>0$, there are infinitely many intersections of $\alpha|_{[0,\varepsilon]}$ and $\gamma$. Let $t$ be such that $\alpha(0)=q=\gamma(t)$. Since both $\alpha$ and $\gamma$ are $\mathcal{C}^{1}$, the fact that $\alpha|_{[0,\varepsilon]}$ intersects $\gamma$ an infinite number of times for every $\varepsilon>0$ implies that $\alpha^{\prime}(0)$ and $\gamma^{\prime}(t)$ are linearly dependent. Since $\gamma^{\prime}(t)\neq 0$ there is some $\lambda\in\mathbb{R}$ with $\alpha^{\prime}(0)=\lambda\cdot\gamma^{\prime}(t)$. If $\lambda\neq 0$, this would be a contradiction to the fact that $\alpha^{\prime}(0)$ is not timelike due to Lemma 4.37. So $\lambda=0$ and hence $\alpha^{\prime}(0)=0$. But $\alpha^{\prime}(0)=d\exp_{tv}(\beta^{\prime}(0))$ Since $\beta^{\prime}(0)\neq 0$ by construction, we conclude that $d\exp_{tv}$ is singular. This means that $q=\exp(tv)$ is conjugate along $\alpha_{v}$, and hence $v\in B$. ∎
4.40 Theorem.
Let $(X,g)$ be a globally hyperbolic spacetime of dimension at least $3$, let $p\in X$, let $L$ be some lightsphere at $p$, let $\{\gamma_{i}\}_{i\in I}$ be some countable family of timelike curves and let $T>0$. If for each future-oriented null geodesic $\alpha$ with $\alpha(0)=p$ and $\alpha^{\prime}(0)\in L$ there is some $t\in(0,T]$ and $i\in I$ such that $\alpha(t)\in\mathrm{im}(\gamma_{i})$, then $(X,g)$ has a compact Cauchy surface and finite fundamental group.
Proof.
By Remark 4.33, we can assume that $(X,g)$ is null-geodesically complete. Further, we can assume without loss of generality that the image of each $\gamma_{i}$ is compact, else subdivide the non-compact $\gamma_{i}$ into countably many overlapping compact timelike curves.
For each $i\in I$ consider the set $A_{i}=\{v\in L\mid\alpha_{v}\text{ intersects }\gamma_{i}\}$. Since the image of $\gamma_{i}$ is compact, each set $A_{i}$ is closed. By assumption $L=\bigcup_{i\in I}A_{i}$. By the Baire category theorem, $L^{\prime}\coloneqq\bigcup_{i\in I}\mathrm{int}(A_{i})$ is dense in $L$. For all $v\in L^{\prime}$, there is some $i\in I$ such that $v\in\mathrm{int}(A_{i})$. By Lemma 4.39 $p$ then has a conjugate point $q$ along $\alpha_{v}|_{(0,T]}$ and therefore $\alpha_{v}(t)\in I^{+}(p)$ for all $t>T$. Application of Lemma 4.35 yields that $(X,g)$ has a compact Cauchy surface.
[…]
$\pi_{1}(S,x)=\#\rho^{-1}(x),$
which is finite since $\tilde{S}$ is compact. ∎
4.41 Example.
Consider $n+1$ dimensional Minkowski spacetime $X=\mathbb{R}^{n}\times\mathbb{R}^{1}_{1}$ and take some countable set of observers $\{\gamma_{i}\}_{i\in I}$ and some $p\in X$. Then the conditions of Theorem 4.40 are not fulfilled. For a null geodesic $\alpha$ with $\alpha(0)=p$ an observer $\gamma$ can intersect at most once with $\alpha$, and so countably many observers can (together) see only countably many light rays. More explicitly, if one takes the set of $\{\gamma_{i}\}_{i\in I}$ to be the set of standard observers $\{t\mapsto(q,t)\}_{q\in\mathbb{Q}^{n}}$ and takes $p=(0,0)$, then all light rays with irrational initial conditions will escape. Theorem 4.40 gives a topological reason for why this has to happen.
The following example shows that the conditions of Theorem 4.40 are sufficient but not necessary for a spacetime to have a compact Cauchy surface with finite fundamental group.
4.42 Example.
Consider the de Sitter spacetime of dimension $n+1$ ($n\geq 2)$, whose Cauchy surfaces are topologically spheres of dimension $n$, so in particular are compact with finite fundamental group. We note that de Sitter space does not meet the conditions of Theorem 4.40. If it did, there would have to be null-geodesics entering into their own chronological future, but this does not happen in de Sitter spacetimes. This makes de Sitter spacetime an example of a globally hyperbolic spacetime with compact Cauchy surfaces with finite fundamental group which is not even very weakly refocusing (and a fortiori not observer-refocusing, strongly refocusing, or refocusing).
4.43 Corollary.
Let $(X,g)$ be a globally hyperbolic spacetime of dimension at least $3$ which is observer-refocusing with respect to a point $p\in X$ and a timelike curve $\gamma$ such that the domain of $\gamma$ is compact. Then $(X,g)$ has a compact Cauchy surface and finite fundamental group.
Proof.
Using similar arguments as in Lemma 4.38 compactness of $\mathrm{im}(\gamma)$ ensures that there is some lightsphere $L$ at $p$ and some $T>0$ such that for all $v\in L$ there is some $t\in(0,T]$ with $\alpha_{v}(t)\in\mathrm{im}(\gamma)$. Now we can use Theorem 4.40. ∎
4.44 Corollary.
Let $(M,h)$ be a $Z^{(x,y)}$ manifold of dimension at least $2$ such that all unit-speed geodesics starting at $x$ reach $y$ in uniformly bounded time. Then $M$ is compact with finite fundamental group.
Proof.
Let $T>0$ be the uniform bound on the time it takes unit-speed geodesics starting at $x$ to reach $y$. Consider the associated spacetime $(X,g)$ to $(M,h)$. Defining $p=(x,0)$ and the standard observer $\gamma:[0,T]\to X$, $\gamma(t)=(y,t)$, the result then follows from Corollary 4.43. ∎
4.45 Corollary.
[…]
Proof.
In this case $M$ is compact and has finite fundamental group. If $\dim(M)=3$, then by the Thurston Elliptization conjecture [34] proved by Perelman [28, 29, 30], $M$ is a quotient of the sphere $S^{3}$ (with standard metric) by a finite group of isometries. The induced metric on $M$ will make it into a $Y^{x}_{2\pi}$ manifold. Similar arguments work for $\dim(M)=2$. ∎
4.46 Remark.
One can carefully check the proof of Lemma 4.39 and Theorem 4.40 to see that if we drop the requirement of a uniform bound $T>0$ from Theorem 4.40 there will still be an open dense subset $A\subset L$ such that all null-geodesics $\alpha$ starting at $p$ with initial velocity in $A$ eventually enter $I^{+}(p)$, but this is not enough to conclude that we have a compact Cauchy surface.
4.47 Remark.
It is unknown whether there are $Z^{x}$ manifolds where the time-of-return of unit-speed geodesics is not uniformly bounded [10, Question 7.71]. Similarly. it is unknown whether there are any globally hyperbolic observer-refocusing spacetimes where the timelike curve cannot be taken to be compact.
5 The Analytic Case
5.48 Theorem.
Let $(X,g)$ be a globally hyperbolic spacetime. Let $p\in X$ and let $L$ be a lightsphere at $p$. Let $v\in L$ such that the geodesic $\alpha_{v}$ has at least $k$ points conjugate to $p$. Then there is some neighborhood $U\subseteq L$ of $v$ such that for each $w\in U$, the null-geodesic $\alpha_{w}$ has at least $k$ conjugate points (counted with multiplicity). Let $\lambda_{1}(w)\leq\lambda_{2}(w)\leq\dots\leq\lambda_{k}(w)$ be the parameter values at which $\alpha_{w}$ hits a conjugate point. Then the functions $\lambda_{j}$ are continuous on $U$ and they are smooth on an open, dense subset of $U$.
Proof.
The existence of the neighborhood $U$ and the continuity of the functions $\lambda_{j}$ are essentially the content of [31, Proposition 4.2]. The fact that each $\lambda_{j}$ is smooth on an open dense set of $U$ follows from [31, Theorem 5.3]: The set $\{\lambda_{j}(v)v\mid v\in U\}$ is a subset of the $T$-conjugate locus as defined in [31]. The regular null $T$-conjugate locus is an open dense subset in the null $T$-conjugate locus which is a smooth submanifold of $\mathcal{N}_{p}X$ transverse to rays in $\mathcal{N}_{p}X$. This implies that for every $j\in\{1,\dots,l\}$ there is an open dense subset $U_{j}\subset U$ such that $\{\lambda_{j}(v)v\mid v\in U_{j}\}$ is a smooth submanifold of $\mathcal{N}_{p}X$ transverse to the rays in $\mathcal{N}_{p}$. By the implicit function theorem the map $v\mapsto\lambda_{j}(v)$ is smooth on $U_{j}$. ∎
5.49 Lemma.
Let $(X,g)$ be a globally hyperbolic spacetime of dimension at least $3$. Let $\gamma$ be some timelike curve in $X$, let $p\in X$ and let $L$ be a lightsphere at $p$. Let $U\subseteq L$ be some open set. Let $f:U\to\mathbb{R}^{+}$ be some lower semi-continuous function such that for all $v\in U$ we have that $\exp(f(v)v)\in\mathrm{im}(\gamma)$ is a point conjugate to p along the geodesic $\alpha_{v}$. Then there is some non-empty open set $V\subseteq U$ on which $\exp(f(v)v)$ is constant.
Proof.
Let the functions $\lambda_{j}$ be as in Theorem 5.48. For $v\in U$ we have that $\exp(f(v)v)$ is a conjugate point along $\alpha_{v}$. Let $k(v)$ denote the minimal $k$ with $\lambda_{k}(v)=f(v)$. We will show that $v\mapsto k(v)$ is lower semi-continuous. Let $v_{n}$ be some sequence in $U$ converging to some $v\in U$ and let $N\in\mathbb{N}$ with $k_{n}\coloneqq k(v_{n})\leq N$ for all $n\in\mathbb{N}$. Then $f(v_{n})=\lambda_{k_{n}}(v_{n})\leq\lambda_{N}(v_{n})$ for all $n\in\mathbb{N}$. By continuity of $\lambda_{N}$ and lower semi-continuity of $f$, we get that $\lambda_{k(v)}(v)=f(v)\leq\lambda_{N}(v)$. Since $k(v)$ is the smallest $k$ with $\lambda_{k}(v)=f(v)$, we must have $k(v)\leq N$. Therefore the function $v\mapsto k(v)$ is lower semi-continuous. For $i\in\mathbb{N}$, let $A_{i}=\{v\in U\mid k(v)\leq i\}$ and note that $U=\bigcup_{i\in\mathbb{N}}A_{i}$. By the Baire category theorem, there is some $i\in\mathbb{N}$ such that $A_{i}$ has non-empty interior. Take $i\in\mathbb{N}$ minimal with this property. Then $V\coloneqq\mathrm{int}(A_{i})\setminus A_{i-1}$ is open. If $\mathrm{int}(A_{i})\setminus A_{i-1}=\emptyset$, then we would have had $\mathrm{int}(A_{i})\subseteq A_{i-1}$ and hence $\mathrm{int}(A_{i})\subseteq\mathrm{int}(A_{i-1})$, contradicting the minimality of $i$. So $V$ is non-empty. For $v\in V$ we have that $f(v)=\lambda_{i}(v)$ and by Theorem 5.48 we can shrink $V$ to some smaller non-empty open set so that $f|_{V}=\lambda_{i}|_{V}$ is smooth. Assume that this $V$ is connected else take a connected component of it instead. Let $s\mapsto v_{s}$ be any $\mathcal{C}^{1}$ curve in $V$ and let us denote $v=f(v_{0})v_{0}$ and $w=\frac{d}{ds}|_{s=0}(f(v_{s})v_{s})$. Then, by Lemma 4.37, $\frac{d}{ds}|_{s=0}\exp(f(v_{s})v_{s})=d\exp_{v}(w)$ is not timelike. But by construction $f(v_{s})\in\mathrm{im}(\gamma)$ for all $s$, so $\frac{d}{ds}f(v_{s})$ is some scalar multiple of a timelike vector. This is only possible if $\frac{d}{ds}|_{s=0}\exp(f(v_{s})v_{s})=d\exp_{v}(w)=0$. Since on $V$ all directional derivatives of $v\mapsto\exp(f(v)v)$ vanish, we conclude that $v\mapsto\exp(f(v)v)$ is constant on $V$. ∎